# Regge trajectories and bridges between them in integrable AdS/CFT

### v2

@@ Abstract
We study the analytic continuation in the spin of the planar spectrum of ABJM theory using the integrability-based Quantum Spectral Curve (QSC) method. Under some minimal assumptions, we classify the analytic properties of the Q functions appearing in the QSC compatible with the spin being non-integer. In this way we find not one - but two distinct possibilities. These two choices correspond to a discrete symmetry of the web of Regge trajectories under a spin shadow transformation: a symmetry which, to the best of our knowledge, had not been discussed before for the Chew-Frautschi plot. Under this symmetry, which we refer to as “twist/co-twist symmetry”, a Regge trajectory is flipped into a “bridge”, which connects leading with sub-leading Regge trajectories. Zigzagging between trajectories and bridges one can reach infinitely many real sheets of the spin Riemann surface without needing to go explicitly around the branch points located in the complex spin plane. Moreover, the symmetry implies the existence of twin points on leading and subleading trajectories, with the same scaling dimensions at any coupling. We discuss how an analogous phenomenon, based on the same mechanism at the level of the QSC, occurs at non-perturbative level in $\mathcal{N}$=4 SYM. This provides a framework to understand some recent independent observations of the symmetry at weak and strong coupling. We present numerical results for Regge trajectories in ABJM theory at finite coupling, in particular we compute exactly the coupling dependence of the position of the leading Regge pole in the correlator of four stress tensors. The shape of this leading trajectory shows a behaviour at weak coupling that strongly resembles the BFKL limit in $\mathcal{N}$=4 SYM.
@@ 1 Introduction
Moreover, something unexpected emerges from our analysis: a new symmetry that pairs different points in the Chew-Frautschi plot of this theory, and which we expect to be there also in $\mathcal{N}$=4 SYM theory!
@@
Figure 1: The plot shows two Regge trajectories of ABJM theory of twists $1$ and $5$, and how they are connected by Bridging trajectories (here two are shown). The Bridges are spin shadows of two other Regge trajectories of twists $4$ and $8$ in this case. This structure underlies the twist/co-twist symmetry: each operator has a shadow sitting on a separate Regge trajectory and related by the exact non-perturbative map $\tau\leftrightarrow\tilde{\tau}$. This is on top of the usual symmetry $\Delta\leftrightarrow d-\Delta$ of the Chew-Frautschi plot. The picture depicts the situation at finite coupling $h=1/10$. The shaded area is the unitarity region, and marked points correspond to physical local states or their shadows. The plot is given in the “top” grading, which means all the shown trajectories interpolate between operators at the top of their supermultiplets.
@@ 1.1 Main results and Twist/Co-twist symmetry
$\mathcal{G}_{\text{Regge}}(u)\simeq\mathcal{G}_{\text{Bridge}}^{T}(u).$ (1)
@@
Combining the properties 1) and 3) above, we arrive at the prediction that the Chew-Frautschi plot of the theory presents a certain symmetry: we can call it twist/co-twist symmetry. [^5] This terminology is inspired by the terminology of conformal spin and co-spin, introduced in JohanH by Johan Henriksson, who noticed what appears to be the same symmetry in 1-loop perturbative results for $\mathcal{N}$=4 SYM. We thank him for exchanges and sharing JohanH with us. Let us call
@@
We are not aware of arguments pointing towards this symmetry in generic CFTs, in fact we believe it is not generic and that, for integrable AdS/CFT models, this statement is new at the non-perturbative level. Perturbatively, what appears to be the same symmetry was noticed by Johan Henriksson at weak coupling in 1-loop studies of $\mathcal{N}$=4 SYM JohanH , by inspecting the results of Henriksson:2017eej for CFT data and noticing that they are symmetric with respect to $\tau\leftrightarrow\tilde{\tau}$. At strong coupling, Julius Julius and Nika Sokolova recently discovered independently a manifestation of the symmetry: they point out that analytic continuation of subleading Regge trajectories is in some cases equivalent to flipping the points with precisely the map (2) (see Julius:2024ewf , in particular Fig. 2). In particular, they found the image of the Konishi operator on a subleading trajectory, consistently with the picture described above.
@@
Notice that our statement of the symmetry leaves something not clarified. Namely, we are saying that a given physical (non-BPS) local operator characterized by $\tau$, $\tilde{\tau}$ with $\tau<\tilde{\tau}$ will have an image on some subleading trajectory with twist $\tilde{\tau}$. However, the number of Regge trajectories increases with respect to the tree-level value of the twist. Therefore, in general there will be multiple trajectories with degenerate tree-level twist approximately $\tilde{\tau}$, and we are not specifying on which one we will find the image of our local operator. Relatedly, we might ask: do all points with integer spin (in steps of 2) present below the unitarity bound on Regge trajectories correspond to spin-flips of local operators according to (2)? Further studies are needed to answer this question and understand the precise pattern with which trajectories are linked by the symmetry, given their tree level quantum numbers. In this paper, we will concentrate on presenting evidence for the general statement of the symmetry written above. [^7] Regarding the more fine-grained question Which trajectory is linked with which by the bridges?, we expect that the answer is hidden in the tree-level form of the Q-functions, which should encode the quantum numbers identifying different trajectories. The transformation rule for Q-functions under the symmetry is explained in this paper. Inspecting the strong coupling data could also be very instructive for decoding the answer to this question, e.g. Julius:2024ewf identifies precisely on which subleading trajectory lives the twist/co-twist image of the Konishi operator.
@@
While we do not know of the possible physical origin of this symmetry, we point out that it is not merely a direct effect of supersymmetry. Notice that supersymmetry implies that Regge trajectories come in groups, shifted by half-integer vectors in the $(\Delta,S)$, plane, replicating the supermultiplet structure, see e.g. Lemos:2021azv . Here, this structure is implicit since the QSC solutions are associated with the whole supermultiplets (disentangling the structure is a matter of susy representation theory). Instead, we are discussing something different: a symmetry that relates points on Regge trajectories which interpolate between different supermultiplets. It would be very interesting, of course, to find an explanation for the symmetry and understand if it is valid also beyond the planar limit.
@@ 2.1 The main ingredients of the QSC
The QSC formalism encodes the spectral data of all non-BPS supermultiplets into the solutions of a particular nonlinear Riemann-Hilbert-type problem. This mathematical problem requires finding some functions of an auxiliary spectral parameter $u$, called Q-functions, based on their analytic properties and the functional relations they satisfy, the so-called QQ relations Reshetikhin:1983vw ; Beisert:2005fw ; Dorey:2006an ; Ferrando:2020vzk ; Ekhammar:2021myw . In short:
@@
The QQ relations are expected to be common to all integrable systems with the same global symmetry. For ABJM theory, these functional relations reflect the superconformal algebra $OSp(4|6)$. On the other hand, the analytic properties are generally model-dependent and will differ, in particular, for integer or non-integer spin.
@@
2.1.1 Quick summary of QQ relations in ABJM theory
@@
As we will see, the functions $\mathbf{P}_{ab}$ and $\mathbf{Q}_{ij}$ have the simplest analytic properties and play a central role in the construction. The essential QQ relations will tell us how to construct $\mathbf{Q}_{ij}$ from $\mathbf{P}_{ab}$. This involves the following two steps:
@@ 2.1.2 Branch cuts and gluing matrix
Figure 2: The cut structure of $\mathbf{P}$ and $\mathbf{Q}$-type Q-functions on their defining Riemann sheet. The position of the branch points depends on the effective coupling $h\equiv h(\lambda)$. The QSC closure axioms describe how to identify the $u$ spectral parameter in the two pictures in such a way that the QQ relations hold.
@@
$\lambda=\frac{\sinh(2\pi h)}{2\pi}\,_{3}F_{1}\left(\frac{1}{2},\frac{1}{2},\frac{1}{2};1,\frac{3}{2};-\sinh^{2}(2\pi h)\right).$ (10)
@@
$\mathbf{P}_{ab}(u)\sim\#\,u^{\mathcal{N}_{ab}},\;\;u\rightarrow+\infty,$ (11)
@@
$\mathbf{Q}_{ij}(u)\sim\#\,u^{\hat{\mathcal{N}}_{ij}},\;\;u\rightarrow+\infty+i0^{+},$ (12)
@@
$Q_{a|j}(u)\sim\#\,u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}},\;\;u\rightarrow+\infty.$ (13)
@@
Figure 3: Using the QQ relations, with the $\mathbf{P}$ functions as input, we can construct two alternative bases for $\mathbf{Q}_{ij}$. Each of these bases is analytic in half of the complex plane and has an infinite series of short cuts in the other half.
@@
$\left.\mathbf{Q}_{ij}(u)\right|_{\text{long cut}}\equiv\mathbf{Q}^{\downarrow}_{ij}(u),\;\;\text{ for }\text{Im}(u)>0,$ (14)
@@
$\left.\mathbf{Q}_{ij}(u)\right|_{\text{long cut}}\equiv\mathcal{G}_{i}^{m}(u)\;{\mathbf{Q}}^{\uparrow}_{mn}(u)\;{\mathcal{G}}_{j}^{n}(u),\;\;\text{ for }\text{Im}(u)<0,$ (15)
@@
Figure 4: The compatibility between the postulated cut structure of Figure 2 and the consequences of the QQ relations in Figure 3 is ensured by gluing the sheets of $\mathbf{Q}^{\downarrow}$ and $\mathbf{Q}^{\uparrow}$. The gluing matrix $\mathcal{G}$ should be a symmetry of the Q-system.
To preserve the integrity of the cut structure, the matrix $\mathcal{G}(u)$ should be free of any cuts or singularities. Additionally, we require $\mathbf{Q}_{ij}$ to satisfy the QQ relations in the lower half plane as well. Consequently, $\mathcal{G}$ must keep the Q-system unchanged, and it is easy to verify that this leads to the following properties:
@@
    which in particular implies that is a periodic matrix under shifts of $u\rightarrow u+2\mathrm{i}$.
@@
Moreover, as we already anticipated, consistency with the analytic properties of Q-functions ensure that
@@ 2.1.3 In-depth discussion of the basis of Q-functions
$Q_{a|i}^{\downarrow}(u)\sim u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}}\left(B_{a|i}^{\downarrow}+\sum_{n=1}^{\infty}\frac{b_{(a|i),n}^{\downarrow}}{u^{n}}\right),\;\;\;u\rightarrow+\infty,$ (18)
which can be analytically extended to the whole upper half plane. Once we fix the leading coefficients [^10] These coefficients should satisfy some constraints which we omit here for brevity but which are easy to determine analysing the large-$u$ expansion of the equations. As an example of an obvious constraint, the coefficients $B_{a|i}$ are constrained by the fact that $\text{det}\left(Q_{a|i}\right)=-1$. $B_{a|i}$, this solution is fixed uniquely if we assume that the powers $\hat{\mathcal{N}}$ are all distinct and non-integer:
@@
For a generic point on a Regge trajectory, this will be the case, but for local operators, $\hat{\mathcal{N}}_{2}-\hat{\mathcal{N}}_{3}$ will be integer since this quantity is identified with (twice) the spin, as we discuss later. When this is the case, we add an extra condition to make the definition of $Q_{a|i}(u)$ unique up to rescalings: if $\hat{\mathcal{N}}_{2}=\hat{\mathcal{N}}_{3}+\ell$ with $\ell\in\mathbb{N}$, we demand that [^11] In this paper, we study only states with $u\leftrightarrow-u$ symmetry, for which this condition is pleonastic, since this coefficient is then automatically zero by the symmetry. In Bombardelli:2018bqz more general states were studied. $b_{(a|2),\ell}=0$ to eliminate the resulting ambiguity in the definition of $Q_{a|2}(u)$, see Bombardelli:2018bqz for full details. Similarly, and with the same subtleties, we define the basis of solutions analytic in the lower half plane as those which have the same asymptotics but for $u\rightarrow-\mathrm{i}\infty$:
$Q_{a|i}^{\uparrow}(u)\sim u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}}\left(B_{a|i}^{\uparrow}+\sum_{n=1}^{\infty}\frac{b_{(a|i),n}^{\uparrow}}{u^{n}}\right).$ (19)
@@
$\Omega_{i}^{m}(u)\kappa_{mn}\Omega_{j}^{n}(u+\mathrm{i})=\kappa_{ij},\;\;\;\;\text{det}(\Omega)=1,$ (22)
@@
$f_{i}^{m}(u)\kappa_{mn}f_{j}^{n}(u+\mathrm{i})=\kappa_{ij},\;\;\;\;\text{det}(f)=1.$ (25)
However, $f_{i}^{j}$, like $\Omega_{i}^{j}$ is not an entire function, but rather is an $2\mathrm{i}$-periodic function with an infinite ladder of branch cuts.
It is known from the TBA derivation inBombardelli:2017vhk that this matrix should have a very peculiar structure:
@@
    where $\mathcal{P}$ is a constant. [^13] This constant may depend on the particular solution of the QSC, and on all parameters except $u$. The 4 functions $\tau_{i}(u)$ should be free of singularities at finite $u$, apart for the branch points.
@@
All in all, the consequence is that the matrix $f_{i}^{j}(u)$ does not have 16 independent elements, but we can reduce its parametrisation to a constant $\mathcal{P}$ and 4 functionally independent degrees of freedom, $\tau_{i}(u)$, subject to (29) and (28). We will use the constraints above in our classification of gluing matrices.

### v3

@@ Abstract
We study the analytic continuation in the spin of the planar spectrum of ABJM theory using the integrability-based Quantum Spectral Curve (QSC) method. Under some minimal assumptions, we classify the analytic properties of the Q-functions appearing in the QSC compatible with the spin being non-integer. In this way we find not one - but two distinct possibilities. While one is related to standard Regge trajectories, we show that the second choice can be used to build bridges which connect leading and subleading Regge trajectories, thus giving a shortcut to reach infinitely many sheets of the spin Riemann surface without going explicitly around the branch points in the complex spin plane. Moreover, the bridges are exact spin reflections of standard Regge trajectories. Together, these results reveal the existence of a hidden symmetry which we call “twist/co-twist symmetry”: every non-BPS local operator has an exact image – living below unitarity on a different Regge trajectory – with the same $\Delta$ and the spin flipped by a Weyl reflection. We discuss how an analogous phenomenon, based on the same mechanism at the level of the QSC, also occurs at non-perturbative level in $\mathcal{N}$$=$$4$ SYM. This provides a framework to understand recent independent observations of the symmetry in this model at weak and strong coupling. We present numerical results for Regge trajectories in planar ABJM theory, in particular we compute exactly the coupling dependence of the position of the leading Regge pole in the correlator of four stress tensors. The shape of this leading trajectory shows a behaviour at weak coupling that strongly resembles the BFKL limit in $\mathcal{N}$$=$$4$ SYM.
@@ 1 Introduction
Moreover, something unexpected emerges from our analysis: a new symmetry that predicts that all the non-BPS, local operator points on the Chew-Frautschi plot have exact spin-flipped images on other trajectories, and which we expect to be there also in $\mathcal{N}$$=$$4$ SYM theory!
@@
Figure 1: The plot shows two Regge trajectories of ABJM theory of twists $1$ and $5$, and how they are connected by Bridging trajectories (here two are shown). The Bridges are spin shadows of two other Regge trajectories of twists $4$ and $8$ in this case. This structure underlies the twist/co-twist symmetry: each operator has a shadow sitting on a separate Regge trajectory and related by the exact non-perturbative map $\tau\leftrightarrow\tilde{\tau}$. This is on top of the usual symmetry $\Delta\leftrightarrow d-\Delta$ of the whole Chew-Frautschi plot Kravchuk:2018htv . The picture depicts some trajectories and bridges at finite coupling $h=0.1$. The shaded area is the unitarity region, and marked points correspond to physical local operators (bullets) or spin-shadows of local operators (circles). The plot is given in the “top” grading, which means all the shown trajectories interpolate between operators at the top of their supermultiplets.
@@ 1.1 Main results and Twist/Co-twist symmetry
$\mathcal{G}_{\text{Regge}}(u)\cong\mathcal{G}_{\text{Bridge}}^{T}(u).$ (1)
@@
Combining the properties 1) and 3) above, we arrive at the prediction that the Chew-Frautschi plot of the theory presents a certain symmetry between pairs of special points: we can call it twist/co-twist symmetry. [^5] This terminology is inspired by the terminology of conformal spin and co-spin, introduced in JohanH by Johan Henriksson, who noticed what appears to be the same symmetry in 1-loop perturbative results for $\mathcal{N}$$=$$4$ SYM. We thank him for exchanges and sharing JohanH with us. Let us call
@@
We are not aware of arguments pointing towards this symmetry in generic CFTs. For integrable AdS/CFT models, to the best of our knowlwedge this statement is new at the non-perturbative level. Perturbatively, what appears to be the same symmetry was noticed by Johan Henriksson at weak coupling in 1-loop studies of $\mathcal{N}$$=$$4$ SYM JohanH , by inspecting the results of Henriksson:2017eej for CFT data and noticing that they are symmetric with respect to $\tau\leftrightarrow\tilde{\tau}$. At strong coupling, Julius Julius and Nika Sokolova recently discovered independently a manifestation of the symmetry: they point out that analytic continuation of subleading Regge trajectories is in some cases equivalent to flipping the points with precisely the map (2) (see Julius:2024ewf , in particular Fig. 2). In particular, they found the image of the Konishi operator on a subleading trajectory, consistent with the picture described above.
@@
Notice that our statement of the symmetry leaves something not clarified. Namely, we are saying that a given physical (non-BPS) local operator characterized by $\tau$, $\tilde{\tau}$ with $\tau<\tilde{\tau}$ will have an image on some subleading trajectory with twist $\tilde{\tau}$. However, the number of Regge trajectories increases with respect to the tree-level value of the twist. Therefore, in general there will be multiple trajectories with degenerate tree-level twist approximately $\tilde{\tau}$, and we are not specifying on which one we will find the image of our local operator. Relatedly, we might ask: since it is not possible that all points with integer spin (in steps of 2) present below the unitarity bound on Regge trajectories correspond to spin-flips of local operators according to (2), what distinguishes the points involved in the symmetry from the ones that are not? Further studies are needed to answer this question and understand the precise pattern with which trajectories are linked by the symmetry, given their tree-level quantum numbers. In this paper, we will concentrate on presenting evidence for the general statement of the symmetry written above. [^7] Regarding the more fine-grained question Which trajectory is linked with which by the bridges?, we expect that the answer is hidden in the tree-level form of the Q-functions, which should encode the quantum numbers identifying different trajectories. The transformation rule for Q-functions under the symmetry is explained in this paper. Inspecting the strong coupling data could also be very instructive for decoding the answer to this question, e.g. Julius:2024ewf identifies precisely on which subleading trajectory lives the twist/co-twist image of the Konishi operator.
@@
While we do not know of the possible physical origin of this symmetry, we point out that it is not merely a direct effect of supersymmetry. Notice that supersymmetry implies that Regge trajectories come in groups, shifted by half-integer vectors in the $(\Delta,S)$, plane, replicating the supermultiplet structure, see e.g. Lemos:2021azv . Here, this structure is implicit since the QSC solutions are associated with the whole supermultiplets (disentangling the structure is a matter of susy representation theory). Instead, we are discussing something different: a symmetry that relates points on Regge trajectories which interpolate between different supermultiplets.
@@
It would be very interesting, of course, to find an explanation for the symmetry and understand whether it is valid also beyond the planar limit. More generally, is it a special feature of integrable holographic theories or perhaps a version of this symmetry might be more general?
@@ 2.1 The main ingredients of the QSC
The QSC formalism encodes the spectral data of all non-BPS supermultiplets into the solutions of a particular nonlinear Riemann-Hilbert-type problem. This mathematical problem requires finding some functions of an auxiliary spectral parameter $u$, called Q-functions, based on their analytic properties and the functional relations they satisfy, the so-called QQ-relations Reshetikhin:1983vw ; Beisert:2005fw ; Dorey:2006an ; Ferrando:2020vzk ; Ekhammar:2021myw . In short:
@@
The QQ-relations are expected to be common to all integrable systems with the same global symmetry. For ABJM theory, these functional relations reflect the superconformal algebra $OSp(4|6)$. On the other hand, the analytic properties are generally model-dependent and will differ, in particular, for integer or non-integer spin.
@@
2.1.1 Quick summary of QQ-relations in ABJM theory
@@
As we will see, the functions $\mathbf{P}_{ab}$ and $\mathbf{Q}_{ij}$ have the simplest analytic properties and play a central role in the construction. The essential QQ-relations will tell us how to construct $\mathbf{Q}_{ij}$ from $\mathbf{P}_{ab}$. This involves the following two steps:
@@ 2.1.2 Branch cuts and gluing matrix
Figure 2: The cut structure of $\mathbf{P}$ and $\mathbf{Q}$-type Q-functions on their defining Riemann sheet. The position of the branch points depends on the effective coupling $h\equiv h(\lambda)$. The QSC closure axioms describe how to identify the $u$ spectral parameter in the two pictures in such a way that the QQ-relations hold.
@@
$\lambda=\frac{\sinh(2\pi h)}{2\pi}\,_{3}F_{2}\left(\frac{1}{2},\frac{1}{2},\frac{1}{2};1,\frac{3}{2};-\sinh^{2}(2\pi h)\right).$ (10)
@@
$\mathbf{P}_{ab}(u)\simeq\#\,u^{\mathcal{N}_{ab}},\quad u\rightarrow+\infty,$ (11)
@@
$\mathbf{Q}_{ij}(u)\simeq\#\,u^{\hat{\mathcal{N}}_{ij}},\quad u\rightarrow+\infty+\mathrm{i}0^{+},$ (12)
@@
$Q_{a|j}(u)\simeq\#\,u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}},\quad u\rightarrow+\infty.$ (13)
@@
Figure 3: Using the QQ-relations, with the $\mathbf{P}_{ab}$ functions as input, we can construct two alternative bases for $\mathbf{Q}_{ij}$. Each of these bases is analytic in half of the complex plane and has an infinite series of short cuts in the other half.
@@
$\left.\mathbf{Q}_{ij}(u)\right|_{\text{long cut}}\equiv\mathbf{Q}^{\downarrow}_{ij}(u),\quad\text{ for }\text{Im}(u)>0,$ (14)
@@
$\left.\mathbf{Q}_{ij}(u)\right|_{\text{long cut}}\equiv\mathcal{G}_{i}^{m}(u)\,{\mathbf{Q}}^{\uparrow}_{mn}(u)\,{\mathcal{G}}_{j}^{n}(u),\quad\text{ for }\text{Im}(u)<0,$ (15)
@@
Figure 4: The compatibility between the postulated cut structure of Figure 2 and the consequences of the QQ-relations in Figure 3 is ensured by gluing the sheets of $\mathbf{Q}^{\downarrow}$ and $\mathbf{Q}^{\uparrow}$. The gluing matrix $\mathcal{G}$ should be a symmetry of the Q-system.
To preserve the integrity of the cut structure, the matrix $\mathcal{G}(u)$ should be free of any cuts or singularities. Additionally, we require $\mathbf{Q}_{ij}$ to satisfy the QQ-relations in the lower half plane as well. Consequently, $\mathcal{G}$ must keep the Q-system unchanged, and it is easy to verify that this leads to the following properties:
@@
    which in particular implies that it is a periodic matrix under shifts of $u\rightarrow u+2\mathrm{i}$.
@@
Moreover, as we already anticipated, consistency with the analytic properties of Q-functions ensures that
@@ 2.1.3 In-depth discussion of the basis of Q-functions
$Q_{a|i}^{\downarrow}(u)\simeq u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}}\left(B_{a|i}^{\downarrow}+\sum_{n\geq 1}\frac{b_{(a|i),n}^{\downarrow}}{u^{n}}\right),\quad u\rightarrow+\infty,$ (18)
which can be analytically extended to the whole upper half plane. Once we fix the leading coefficients [^10] These coefficients should satisfy some constraints which we report in Appendix B. As an example of an obvious constraint, the coefficients $B_{a|i}^{\downarrow}$ should be compatible with the fact that $\text{det}\left(Q_{a|i}^{\downarrow}\right)=-1$. $B_{a|i}^{\downarrow}$, this solution is fixed uniquely if we assume that the powers $\hat{\mathcal{N}}$ are all distinct and non-integer:
@@
For a generic point on a Regge trajectory, this will be the case, but for local operators, $\hat{\mathcal{N}}_{2}-\hat{\mathcal{N}}_{3}$ will be integer since this quantity is identified with (twice) the spin, as we discuss later. When this is the case, we add an extra condition to make the definition of $Q_{a|i}^{\downarrow}(u)$ unique up to rescalings: if $\hat{\mathcal{N}}_{2}=\hat{\mathcal{N}}_{3}+\ell$ with $\ell\in\mathbb{N}$, we demand that [^11] In this paper, we study only states with $u\leftrightarrow-u$ symmetry, for which this condition is pleonastic, since this coefficient is then automatically zero by the symmetry. In Bombardelli:2018bqz more general states were studied. $b_{(a|2),\ell}^{\downarrow}=0$ to eliminate the resulting ambiguity in the definition of $Q_{a|2}^{\downarrow}(u)$, see Bombardelli:2018bqz for full details. Similarly, and with the same subtleties, we define the basis of solutions analytic in the lower half plane as those which have the same asymptotics but for $u\rightarrow-\mathrm{i}\infty$:
$Q_{a|i}^{\uparrow}(u)\simeq u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}}\left(B_{a|i}^{\uparrow}+\sum_{n\geq 1}\frac{b_{(a|i),n}^{\uparrow}}{u^{n}}\right).$ (19)
@@
$\Omega_{i}^{m}(u)\kappa_{mn}\Omega_{j}^{n}(u+\mathrm{i})=\kappa_{ij},\quad\text{det}(\Omega)=1,$ (22)
@@
$f_{i}^{m}(u)\kappa_{mn}f_{j}^{n}(u+\mathrm{i})=\kappa_{ij},\quad\text{det}(f)=1.$ (25)
However, $f_{i}^{j}$, like $\Omega_{i}^{j}$ is not an entire function, but rather is an $2\mathrm{i}$-periodic function with an infinite ladder of branch cuts. It is known from the TBA derivation in Bombardelli:2017vhk that this matrix should have a very peculiar structure:
@@
    where $\mathcal{P}$ is a constant. [^13] This constant may depend on the particular solution of the QSC, and on all parameters except $u$. The 4 functions $\tau_{i}(u)$ should be free of singularities at finite $u$, apart from the branch points.
@@
All in all, the consequence is that the matrix $f_{i}^{j}(u)$ does not have 16 independent elements, but we can reduce its parametrisation to a constant $\mathcal{P}$ and four functionally independent degrees of freedom, $\tau_{i}(u)$, subject to (29) and (28). We will use the constraints above in our classification of gluing matrices.